If Y = 1/X , write down the value of Y, correct to 2 decimal places, when X = 8
Answer by umail08(87)

If y = 1/x … when x = 8 then
y = 1/x
y = 1/8
y = 0.125
y = 0.13 …. Rounded to 2 decimals places